OpenStudy (anonymous):

A line passes through (6,-1) with a slope of -2. What is the Point-Slope form of the equation of this line?

OpenStudy (aravindg):

point slope form is y-y1=m(x-x1)

OpenStudy (aravindg):

where (x1,y1) is the point and m is slope

OpenStudy (anonymous):

A) -2(y+1)=x-6 B) y-1=-2(x+6) C) y=-2x+1 D) y+1=-2 (x-6) E) None of the above

OpenStudy (aravindg):

please do that using the hints i gave,,we do not give direct answers here .we help u get to it !